Griffin's Goat Farm, Inc., has sales of $644,000, costs of $285,000, depreciation expense of $53,000, interest expense of $20,000, and a tax rate of 21 percent. What is the net income for this firm?
Answer #1

Calculation of net income:

Earning before interest and tax= sales-cost-depreciation

Earning before interest and tax= 644000-285000-53000

Earning before interest and tax= 306000

Earning before tax= Earning before interest and tax- interest

Earning before tax= 306000-20000= 286000

Earning after tax= 286000*(1-0.21)= 286000*0.79= 225940

Net income= 225940